If sin(pi/2 -x)=0.35, what's cos x?
A).35
B).43
C).45
D).53

Answer by ikleyn(52943)

.
If sin(pi/2 -x)=0.35, what's cos x?

A) .35     <<<---===  sin(pi/2-x) = cos(x).     MEMORIZE it   ! ! !

B) .43

C). 45

D) .53